let f(x)= 6+x and g(x)= x^2-x+1, find (f-g)(x)
Answer by stanbon(75887)

let f(x)= 6+x and g(x)= x^2-x+1, find (f-g)(x)
-----
(f-g)(x) = f(x)-g(x) = (6+x)-(x^2-x+1)
---
= -x^2+2x+5
==============
Cheers,
Stan H.